# Strain-Enhanced Altermagnetism in Ca3Ru2O7

### v1

@@ 3.3 Tuning AFM to AM phases
A direct comparison between relaxed structure and the non-relaxed structure allows us to disentangle the effects of strain-induced delocalization/localization from those arising purely due to octahedral distortions. In particular, the density of states (DOS) in Fig. S8 confirms that compressive strain broader states, reducing AM features. In contrast, tensile strain narrows the bandwidth, reinforcing the AM character. Moreover, the band structures shown in Fig. S10 illustrate how freezing the atomic positions primarily alters the gap structure rather than the AM spin splitting itself, reinforcing the observation that the observed strain-dependent AM behavior is primarily dictated by electronic exchange interactions, with octahedral distortions acting as a secondary fine-tuning mechanism.
Furthermore, an interesting behavior can be observed in the evolution of the distortion angle () as a function of strain (see Fig. 6(d)). In configuration A, increases monotonically from to strain. However, in configuration B, exhibits a more complex response, increasing under compressive strain () but displaying a non-monotonic behavior under tensile strain. This difference originates from the distinct in-plane magnetic configurations of the two phases. Compressive strain generally enhances electronic delocalization, which tends to weaken magnetic exchange interactions and favor a more symmetric crystal structure. In the case of conf. A, which is metallic (see, Fig. 3a)), the effect of in-plane strain gradually increases delocalization, reducing structural distortions and stabilizing a more symmetric octahedral environment. Conversely, electronic states are more localized in configuration B (see, Fig. 3b)), where the in-plane magnetic alignment is AFM. As a result, under compression, the system seeks to accommodate strain by increasing octahedral distortions (), effectively enhancing localization to maintain magnetic interactions, resulting in a more distorted structure than configuration A. Under tensile strain, however, competing effects emerge: while in-plane expansion reduces the need for further distortions, the slight contraction along the -axis counterbalances this effect, leading to either an increase or decrease in , depending on the precise strain conditions.
The strain-dependent behavior of is closely tied to the observed changes in the energy profile in Fig. 6(a), where configuration B becomes energetically favorable at compressive strain values above . The stabilization of the AM phase coincides with an increase in , suggesting that octahedral distortions play a crucial role in the strain-driven AFM-to-AM transition.
4 Final Remarks
In this work, we explore the emergence of altermagnetic states in the correlated material Ca3Ru2O7. While its ground state lacks AM features, alternative magnetic configurations enable their realization. AM arises in spin arrangements connected through rotational and translational symmetries. In particular, configuration B, which alternates spin orientation both in-plane and between layers, exhibits spin connectivity via rotations along the -axis and translations in the - plane. This configuration hosts a P-2 -wave altermagnetic state, which is orbital-selective, with AM primarily arising from the - orbitals. When the spin-orbit coupling is included, with the Néel vector oriented along the -axis, time-reversal symmetry breaking leads to weak ferromagnetism along the -axis due to the staggered Dzyaloshinskii-Moriya interaction [7, 15]. While configuration B has not been experimentally observed in pristine Ca3Ru2O7, it has been reported in doped systems, such as Ca3(TixRu1-x)2O7, suggesting that this system could host an AM ground state under suitable conditions. Additionally, our results indicate that strain plays a crucial role in tuning AM properties. Tensile strain along the -axis enhances AM band splitting by approximately (for strain), whereas compressive strain suppresses it by . Moreover, we predict a strain-driven phase transition from an AFM to an AM state (configuration A to B) under compressive strain exceeding . While previous experiments have manipulated magnetic anisotropy using uniaxial strain (), the impact of biaxial strain remains largely unexplored.
Strain-tuned AM phases could significantly expand the known quantum phenomena in the Can+1RunO3n+1 family, which already exhibit strain-driven transitions, such as Mott phase suppression in Ca2RuO4 [17], the emergence of a Kondo effect, and a metal-to-semiconductor transition in CaRuO3 ( strain) [42]. Our findings provide a fundamental step toward uncovering possible AM states in the Ca3Ru2O7 family and pave the way for future theoretical and experimental investigations of biaxial strain as a tool for engineering AM phases in correlated oxides.
@@ Acknowledgments
A.L. thanks the Technische Universität Dresden Professor Fellowship Program. C. A. acknowledges the useful discussion with G. Cuono and A. Fakhredine about altermagnetism in Ca2RuO4. C. A. is supported by the Foundation for Polish Science project “MagTop” no. FENG.02.01-IP.05-0028/23 co-financed by the European Union from the funds of Priority 2 of the European Funds for a Smart Economy Program 2021–2027 (FENG). J.W.G. acknowledges financial support from Chilean ANID-FONDECY grants N. 1220700 and 1221301. T.B. would like to thank the German Science Foundation (DFG) for supporting this work via the German-Israeli Project Cooperation with the Grant No. HE 3543/42-1.

### v2

@@ Discussion
A direct comparison between the relaxed structures and the non-relaxed structures allows us to disentangle the effects of strain-induced delocalization/localization from those arising purely from octahedral distortions. In particular, the density of states in Fig. S5 confirms that compressive strain broader states, reducing AM features. In contrast, tensile strain narrows the bandwidth, reinforcing the AM character. Moreover, the band structures shown in Fig. S7 illustrate how freezing the atomic positions primarily alters the gap structure rather than the AM spin splitting itself, reinforcing the observation that the observed strain-dependent AM behavior is primarily dictated by electronic exchange interactions, with octahedral distortions acting as a secondary fine-tuning mechanism.
Furthermore, Fig. 8(d) reveals how the RuO octahedral distortion angle responds to biaxial strain in each magnetic configuration. In configuration A, increases smoothly from to strain. Configuration B shows a more intricate pattern: under compressive strain (), increases as the lattice compresses in the plane. This response reflects stronger electronic localization in the antiferromagnetic arrangement (see Fig. 3(b)), which preserves magnetic exchange by enhancing octahedral distortion. Under tensile strain, in-plane expansion tends to relax distortions while the slight contraction along the axis reinforces them. The net change in depends on the balance between these two opposing effects.
The evolution of in configuration B correlates directly with the energy profile in Fig. 8(a). Above compressive strain, configuration B becomes lower in energy than configuration A and increases sharply. This coincidence demonstrates that enhanced octahedral distortion drives the AFM-to-AM transition. Although such strain levels are challenging to achieve in bulk crystals, epitaxial oxide films can routinely accommodate comparable lattice mismatches, providing a practical platform to realize and probe distortion-driven magnetic and electronic phase transitions in layered perovskites [45, 46, 47, 48].
This work investigates the emergence of altermagnetic states in the correlated oxide Ca3Ru2O7. While its ground state is not altermagnetic, symmetry-allowed AM order arises in alternative spin configurations. In the non-relativistic case, Configuration B, characterized by a Néel-type spin pattern, hosts a P-2 -wave altermagnetic state.
Including spin-orbit coupling with the Néel vector along induces additional symmetry breaking, allowing a -wave component in the spin texture. This hybridization between - and -wave features may result from the absence of inversion symmetry and the presence of nonsymmorphic operations. A complete understanding of the -wave origin will require further analysis based on spin space group symmetries under non-collinear magnetism.
Although configuration B has not been reported in pristine Ca3Ru2O7, it appears in doped compounds such as Ca3(TixRu1-x)2O7, suggesting that suitable conditions can stabilize the AM state. Our results further demonstrate that biaxial strain acts as an effective tuning parameter for the altermagnetic band splitting: tensile strain (or equivalently, compression along the -axis) enhances the splitting by approximately at strain, whereas compressive strain reduces it by about . We also predict a strain-induced transition from the AFM ground state to the AM state beyond 2% compressive strain. While prior studies have used uniaxial strain to manipulate magnetic anisotropy [26], the effects of biaxial strain remain largely unexplored.
Inducing altermagnetism in Ca3Ru2O7 would create a rare noncentrosymmetric altermagnetic material and provide a platform to explore the interplay between odd- and even-wave spin-momentum locking. Identifying such phases would broaden the spectrum of quantum phenomena known in the Can+1RunO3n+1 family, which already exhibits strain-tunable transitions, including Mott suppression in Ca2RuO4 ( strain)[18], the emergence of a Kondo effect, and a metal-to-semiconductor transition in CaRuO3 ( strain)[45]. Our results highlight biaxial strain as a promising route to control altermagnetic phases in Can+1RunO3n+1-related materials, aligning with recent proposals to manipulate antiferromagnetic-altermagnetic phase transitions via strain [49, 50]. These findings motivate further exploration of symmetry-driven spin textures in systems with strong spin-orbit coupling.
Note: Recent studies have shown that SOC can induce interesting phenomena such as polarization switching in Ca3Mn2O7, driven by magnetic anisotropy [43]. Given its structural similarity to Ca3Ru2O7, our results suggest that Ca3Mn2O7 may also host mixed-symmetry altermagnetic order under SOC. Further exploration of this system could offer new pathways for tuning altermagnetic behavior via SOC.
Methods
We perform a theoretical analysis using the density functional theory with and without spin-orbit coupling. We use the plane-wave pseudopotential method implemented in the Vienna ab-initio simulation package (VASP)[51] within the generalized gradient approximations (GGA), and it gives structural properties closer to the experimental values. The electronic valence considered are: Ru: 54 and O: 22, and Ca: 334. We use a plane-wave energy cutoff of 650 eV and set a Regular Monkhorst-Pack grid of 775 to perform the atomic relaxation and 11117 to perform the self-consistent calculation. We use a fine k-grid of 14147 within the tetrahedron method for the density of states. We perform the structural optimization of the unit cell until a force convergence threshold of at least 10-3 eV/Å per atom.
To account for electronic correlations in the Ru -orbitals, we use the Dudarev approach [52] with an effective Hubbard parameter eV, previously shown to yield accurate structural and electronic properties for CRO [36]. We use a Cartesian coordinate system with crystallographic axes aligned as , , and .
Data availability
All data supporting the findings of this study are included in the article and its Supplementary Information.
Code Availability
Density functional theory calculations were performed using the Vienna ab initio Simulation Package (VASP), accessed under an institutional license through the high-performance computer at the NHR Center of TU Dresden and on the high-performance computers Noctua 2 at the NHR Center PC2. Custom Python scripts developed for post-processing and analysis are available from the corresponding author upon reasonable request.
@@ Acknowledgments
A.L. thanks the Technische Universität Dresden Professor Fellowship Program. C. A. acknowledges the useful discussion with G. Cuono and A. Fakhredine about altermagnetism in Ca2RuO4. C. A. is supported by the Foundation for Polish Science project “MagTop” no. FENG.02.01-IP.05-0028/23 co-financed by the European Union from the funds of Priority 2 of the European Funds for a Smart Economy Program 2021-2027 (FENG). J.W.G. acknowledges financial support from Chilean ANID-FONDECY grants N. 1220700 and 1221301. T.B. would like to thank the German Science Foundation (DFG) for supporting this work via the German-Israeli Project Cooperation with the Grant No. HE 3543/42-1. The authors gratefully acknowledge the computing time made available to them on the high-performance computer at the NHR Center of TU Dresden and on the high-performance computers Noctua 2 at the NHR Center PC2.
Author Contributions
A.L. and J.G. performed the calculations and contributed to the organization of the article. A.L. and C.A. held foundational discussions about the study. A.L., J.G., and C.A. carried out the symmetry analysis. A.L. and T.B. performed the electronic structure analysis. A.L., J.G., and C.A. wrote the initial draft. All authors contributed to the discussion of the results and collaborated in editing and shaping the final version of the manuscript.
Competing interests
The authors declare no competing interests.